OpenStudy (anonymous):

11. 3x+2y=1 4x+ 6y= 7

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy :) I'm assuming you would like to solve for x and y from those two equations would you like to use substitution or elimination method?

OpenStudy (anonymous):

Solve for y or x

OpenStudy (anonymous):

Using the system of the elimination method.

jigglypuff314 (jigglypuff314):

ok :) then first multiply through the first equation by 3 do you know how to do that? what would you get?

OpenStudy (anonymous):

9x+6y=3 for the first equation

jigglypuff314 (jigglypuff314):

good! :) then we can line it up like 9x + 6y = 3 4x + 6y = 7 and subtract

OpenStudy (anonymous):

x=5

jigglypuff314 (jigglypuff314):

hmm? 9x - 4x = 5x 6y - 6y = 0 3 - 7 = -4 so we get 5x = -4 solve for x from there

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have one more question on a different problem can you help me

jigglypuff314 (jigglypuff314):

you're welcome :) then once you get the x, you can plug it back into one of the original equations and solve for y

jigglypuff314 (jigglypuff314):

of course, ask away ^_^

OpenStudy (anonymous):

4x+6y=8 2x+3y=3 using elimation

jigglypuff314 (jigglypuff314):

multiply the second equation by 2 subtract then know the special cases 0=0 is a true statement and means that the two equations are over lapping and share an infinite solutions 0=(any other real number) is a false statement and means that the two equations are parallel and so have no solutions

OpenStudy (anonymous):

ok thank you i have having trouble seeing if i was right. you saved my day thanks. Have a great night. ^_-

jigglypuff314 (jigglypuff314):

Glad I could help ^_^ <3